answersLogoWhite

0

🎒

Statistics

Statistics deals with collecting, organizing, and interpreting numerical data. An important aspect of statistics is the analysis of population characteristics inferred from sampling.

36,755 Questions

What is the record for one woman having the most children?

The world record for having the most number of children officially recorded is 69 by the first of two wives of Feodor Vassilyev (1707-1782), a peasant from Shuya, 150 miles east of Moscow. In 27 confinements, she gave birth to 16 pairs of twins, seven sets of triplets and four sets of quadruplets. The children were born between 1725-1765.

Trying to find what the phrase where the bear st in the buckwheat means?

Where the bear s**t in the buckwheat, alludes to something obvious, as in, "You don't have to tell me where the bear s**t in the buckwheat." Less confused would be to say "that a bear s**ts in the buckwheat (nature/forest). I'm not sure, but it's probably related to "a bear s**ts in the woods", and just another way to say the same thing.

What if your instructor asks you to determine the number of organisms in a water sample observation of your dilution plates reveals?

I think you would eliminate plate counts that are not between 30-300 colonies. <30, because its too few to count, >300 too numerous to count.

How many four-digit numbers can you make by arranging the numbers 1 2 8 9?

If each of the numbers 1, 2, 8, and 9, each of which is also a digit, must be used exactly once, the answer is 4 X 3 X 2 = 24. Reasoning: Any one of the four digits may be taken first; any one of the remaining three may be taken second, and either of the remaining two may be taken next, leaving only one choice for the final digit.

A distribution statement does what?

identifies who is authorized receipt of the to

How do you find the probability of drawing 5 red cards in 5 card poker hand?

A quick way is to simply count all the possible hands with red cards and compare to the total number of hands. There are 52 total cards and 26 red cards. We can use the binomial coefficient for this, otherwise known as the "choose" operator.

A = Total Hands: (52c5) = 2598960
B = Red Hands: (26c5) = 65780
C = Chance of red hand = A / B = 2.53%

Alternately you can use a permutation solution, for each card, in order, what is the chance to get a red card. There are 26 of 52 red cards on the first card, so 26/52. Then continue:

26/52 * 25/51 * 24/50 * 23/49 * 22/48 = 2.53%

If you play with the form of the equations, and understand a bit about combinations, you'll see that this second form is very closely related to the expanded version of the first form. This two methods have advantages in different situations. Here they are about the same.

What is the mean of 7 8 20 6 9 11 10 8 9 8?

The mean of these, or any group of ten numbers, is 1/10th of their sum.

The letter P in combination with what number indicates a brain dead patient?

P6 - Declared brain-dead patient whose organs are being removed for donor purposes

Is the mean and median mode and range always part of a data set?

yes it is always

* * * * *

That is pure rubbish. The mean and median score for a throw of a die are 3.5 But there are not many dice with a face value of 3.5!

The mode must be - always.

The range often is not.

If the height of adult males ranges in height from a minimum of 1.5 metres to a maximum of 2.0 metres, then the range is 0.5 metres. However, the fact that the minimum is 1.5 m implies there is no male with a height of 0.5 m. So, the range is not part of the data set.

How many 3 digit combinations when using 10 numbers once?

10 x 9 x 8 = 720 different arrangements

If the sequence doesn't matter, then there are (720/6) = 120 different bunches

What are the reasons in point form girls blush dark red with only you if you are a Muslim guy?

They all probably like you and find you attractive, this is common. I do not believe ethnicity has to do anything with it. I noticed you have asked this question several times also.

What are the possible combinations of rolling two dice to get 11?

11 = 6+5 is the only solution, so there are two combinations

first dice : 6, second dice : 5

first dice : 5, second dice : 6

What are 5 numbers digits are the same between 0 and 60?

If I understand the question correctly,

the digits 1, 2, 3, 4 and 5 appear the same number of times.

How many digits are there in an alien a number?

Alien registration number should be 9 digits long with no alphabetical characters.